# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v6

@@ F.3 Technical Lemmas for Lemma 1
$\mathbf{w}_{t+1}=\arg\min_{\mathbf{w}\in\mathcal{V}}\eta\ell_{t}(\mathbf{w})+D_{\psi}(\mathbf{w},\mathbf{w}_{t}),$
[…]
@@
Lemma F.5 (Lemma 15 of Zhang and Sugiyama 52).
@@
Lemma F.6 (Lemma 1 of Zhang and Sugiyama 52).
@@
Lemma F.7 (Lemma 17 of Zhang and Sugiyama 52).
Let $\mathbf{z}\in\mathbb{R}^{K}$ be a $K$-dimensional vector. Let $\ell(\mathbf{z},\mathbf{y})=\sum_{k=0}^{K}\mathbf{1}\{y_{i}=1\}\cdot\log\left(\frac{1}{[\bm{\sigma}(\mathbf{z})]_{k}}\right)$, where $\mathbf{y}=[y_{0},\dots,y_{K}]^{\top}\in\mathbb{R}^{K+1}$, and the softmax function $\bm{\sigma}(\mathbf{z}):\mathbb{R}^{K}\rightarrow\mathbb{R}^{K}$ is defined as $[\bm{\sigma}(\mathbf{z})]_{i}=\frac{\exp([\mathbf{z}]_{i})}{v_{0}+\sum_{k=1}^{K}\exp([\mathbf{z}]_{k})}$ for all $i\in[K]$, and $[\bm{\sigma}(\mathbf{z})]_{0}=\frac{v_{0}}{v_{0}+\sum_{k=1}^{K}\exp([\mathbf{z}]_{k})}$. Define $\mathbf{z}^{\mu}:=\bm{\sigma}^{+}\left(\operatorname{smooth}_{\mu}(\bm{\sigma}(\mathbf{z}))\right)$, where $\operatorname{smooth}_{\mu}(\mathbf{q})=(1-\mu)\mathbf{q}+\mu\mathbf{1}/(K+1)$. Then, for $\mu\in[0,1/2]$, we have
@@
Lemma F.8 (Lemma 18 of Zhang and Sugiyama 52).
Let $L_{t}(\mathbf{w})=\ell_{t}(\mathbf{w})+\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{t}\|_{H_{t}}^{2}$. Assume that $\ell_{t}$ is a $M$-self-concordant-like function. Then, for any $\mathbf{w},\mathbf{w}_{t}\in\mathcal{W}$, the quadratic approximation $\tilde{L}_{t}(\mathbf{w})=L_{t}(\mathbf{w}_{t+1})+\langle\nabla L_{t}(\mathbf{w}_{t+1}),\mathbf{w}-\mathbf{w}_{t+1}\rangle+\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{t+1}\|_{H_{t}}^{2}$ satisfies
@@ Appendix G Proofs of Theorem 3
In this section, we provide the proof of Theorem 3. In addition to the adversarial construction presented in Section D.1, we construct the adversarial non-uniform rewards.
@@ G.1 Adversarial Rewards Construction
Under the adversarial construction in Section D.1, we observe that there are $K$ identical context vectors, invariant across rounds $t$. Therefore, in total, there are $N=K\cdot\binom{d}{d/4}$ items. Let the rewards be also time-invariant. Given $\mathbf{w}_{V}$, we define a unique item $i^{\star}\in[N]$ as an item that maximizes $x_{i}^{\top}\mathbf{w}_{V}$, i.e., $x_{i^{\star}}=x_{V}$, and has a reward of $1$, i.e., $r_{i^{\star}}=1$. Then, we construct the non-uniform rewards as follows:
$r_{i}=\begin{cases}1,&\text{for }i=i^{\star}\\ \gamma,&\text{for }i\neq i^{\star},\end{cases}$ (G.1)
where we define $\gamma$ as
$\gamma=\min_{S\in\mathcal{S}}\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}=\frac{1}{v_{0}+1}.$
Note that $\gamma<1$.
@@ G.2 Main Proof of Theorem 3
Given the rewards construction as (G.1), any reward in the optimal assortment $S_{t}^{\star}$ is larger than the expected revenues.
@@
Let $R(S^{\star},\mathbf{w}_{V})=\frac{\sum_{i\in S^{\star}}\exp(x_{i}^{\top}\mathbf{w}_{V})r_{i}}{v_{0}+\sum_{j\in S^{\star}}\exp(x_{i}^{\top}\mathbf{w}_{V})}$. Then, we have
$r_{i}\geq R(S^{\star},\mathbf{w}_{V}),\quad\forall i\in S^{\star}.$
Lemma G.1 implies that $S^{\star}$ contains only one item $i^{\star}$. This is because if $S^{\star}=\{x_{i^{\star}}\}$, adding any item $i\neq i^{\star}$ to the assortment results in lower expected revenue, since $r_{i}=\gamma\leq R(S^{\star}=\{x_{i^{\star}}\},\mathbf{w}_{v})$.
Furthermore, we can bound the expected revenue for any assortment as follows:
@@
Under the same parameters and context vectors as those in Section D, if the rewards are constructed according to Equation (G.1), for any $S\in\mathcal{S}$, we have
$R(S,\mathbf{w}_{V})\leq\frac{\max_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\max_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}.$
Let $x_{U_{1}},\dots,x_{U_{L}}$ be the distinct feature vectors contained in assortments $S_{t}$ with $U_{1},\dots,U_{L}\in\mathcal{V}_{d/4}$. Let $U^{\star}$ be the subset among $U_{1},\dots,U_{L}$ that maximizes $x_{U}^{\top}\mathbf{w}_{V}$, i.e., $U^{\star}\in\mathop{\mathrm{argmax}}_{U\in\{U_{1},\dots,U_{L}\}}x_{U}^{\top}\mathbf{w}_{V}$, where $\mathbf{w}_{V}$ is the underlying parameter. For simplicity, we denote $\tilde{U}_{t}$ as the unique $U^{\star}\in\mathcal{V}_{d/4}$ in $S_{t}$. Then, we have
$\sum_{t=1}^{T}R(S^{\star},\mathbf{w}_{V})-R(S_{t},\mathbf{w}_{V})$ $=\sum_{t=1}^{T}\frac{\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{V}^{\top}\mathbf{w}_{V})}-R(S_{t},\mathbf{w}_{V})$
$\geq\sum_{t=1}^{T}\frac{\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{V}^{\top}\mathbf{w}_{V})}-\frac{\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}$
$=\sum_{t=1}^{T}\frac{\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{V}^{\top}\mathbf{w}_{V})}-\frac{\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})},$
where the first equality holds becauset $S^{\star}$ contains only one item $i^{\star}$ by Lemma G.1 (and recall that $x_{i^{\star}}=x_{V}$), and the inequality holds by Lemma G.2. Hence, the problem is not easier than solving the MNL bandit problems with the assortment size $1$, i.e., $K=1$. By putting $K=1$ and $v_{0}=\Theta(1)$ in Theorem 1, we derive that
$\sup_{\mathbf{w}}\mathbb{E}_{\mathbf{w}}^{\pi}\left[\textup{{Reg}}_{T}(\mathbf{w})\right]$ $\geq\frac{1}{|\mathcal{V}_{d/4}|}\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{\mathbf{w}_{V}}^{\pi}\sum_{t=1}^{T}R(S^{\star},\mathbf{w}_{V})-R(S_{t},\mathbf{w}_{V})$
$\geq\frac{1}{|\mathcal{V}_{d/4}|}\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{\mathbf{w}_{V}}^{\pi}\sum_{t=1}^{T}\frac{\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{V}^{\top}\mathbf{w}_{V})}-\frac{\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}$
$=\Omega\left(d\sqrt{T}\right).$
@@ G.3 Proofs of Lemmas for Theorem 3
We prove by contradiction. Assume that there exists $i\in S^{\star}$ such that $r_{i}<R(S^{\star},\mathbf{w}_{V})$. Then, removing the item $i$ from the assortment $S^{\star}$ yields higher expected revenue. This contradicts the optimality of $S^{\star}$. Thus, we have
$r_{i}\geq R(S^{\star},\mathbf{w}_{V}),\quad\forall i\in S^{\star}.$
This concludes the proof. ∎
@@
Case 1. $i^{\star}\in S$.
@@
$\gamma=\min_{S\in\mathcal{S}}\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}.$ (G.2)

### v9

@@ F.3 Technical Lemmas for Lemma 1
$\mathbf{w}_{t+1}=\arg\min_{\mathbf{w}\in\mathcal{V}}\eta_{t}\ell_{t}(\mathbf{w})+D_{\psi}(\mathbf{w},\mathbf{w}_{t}),$
where $\mathcal{V}\subseteq\mathcal{W}\subseteq\mathbb{R}^{d}$ is a non-empty convex set and $D_{\psi}(\mathbf{w}_{1},\mathbf{w}_{2})=\psi(\mathbf{w}_{1})-\psi(\mathbf{w}_{2})-\langle\nabla\psi(\mathbf{w}_{2}),\mathbf{w}_{1}-\mathbf{w}_{2}\rangle$ is the Bregman Divergence w.r.t. a strictly convex and continuously differentiable function $\psi:\mathcal{W}\rightarrow\mathbb{R}$. Further supposing $\psi(\mathbf{w})$ is $1$-strongly convex w.r.t. a certain norm $\|\cdot\|$ in $\mathcal{W}$, then there exists a $\mathbf{g}^{\prime}_{t}\in\partial\ell_{t}(\mathbf{w}_{t+1})$ such that
@@
Lemma F.5 (Lemma 15 of Zhang and Sugiyama 51).
@@
Lemma F.6 (Lemma 1 of Zhang and Sugiyama 51).
@@
Lemma F.7 (Lemma 17 of Zhang and Sugiyama 51).
Let $\mathbf{z}\in\mathbb{R}^{K}$ be a $K$-dimensional vector. Let $\ell(\mathbf{z},\mathbf{y})=\sum_{k=0}^{K}\mathbf{1}\{y_{i}=1\}\cdot\log\left(\frac{1}{[\bm{\sigma}(\mathbf{z})]_{k}}\right)$, where $\mathbf{y}=[y_{0},\dots,y_{K}]^{\top}\in\mathbb{R}^{K+1}$, and the softmax function $\bm{\sigma}(\mathbf{z}):\mathbb{R}^{K}\rightarrow\mathbb{R}^{K}$ is defined as $\bm{\sigma}(\mathbf{z})]_{i}=\frac{\exp([\mathbf{z}]_{i})}{v_{0}+\sum_{k=1}^{K}\exp([\mathbf{z}]_{k})}$ for all $i\in[K]$, and $\bm{\sigma}(\mathbf{z})]_{0}=\frac{v_{0}}{v_{0}+\sum_{k=1}^{K}\exp([\mathbf{z}]_{k})}$. Define $\mathbf{z}^{\mu}:=\bm{\sigma}^{+}\left(\operatorname{smooth}_{\mu}(\bm{\sigma}(\mathbf{z}))\right)$, where $\operatorname{smooth}_{\mu}(\mathbf{q})=(1-\mu)\mathbf{q}+\mu\mathbf{1}/(K+1)$. Then, for $\mu\in[0,1/2]$, we have
@@
Lemma F.8 (Lemma 18 of Zhang and Sugiyama 51).
Let $L_{t}(\mathbf{w})=\ell_{t}(\mathbf{w})+\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{t}\|_{H_{t}}^{2}$. Assume that $\ell_{t}$ is a $M$-self-concordant-like function. Then, for any $\mathbf{w},\mathbf{w}_{t}\in\mathcal{W}$, the quadratic approximation $\tilde{L}_{t}(\mathbf{w})=L_{t}(\mathbf{w}_{t+1})+\langle\nabla^{2}L_{t}(\mathbf{w}_{t+1}),\mathbf{w}-\mathbf{w}_{t+1}\rangle+\frac{1}{2c}\|\mathbf{w}-\mathbf{w}_{t+1}\|_{H_{t}}^{2}$ satisfies
@@
Lemma F.9.
Let $\mathbf{w}_{t+1}=\mathop{\mathrm{argmin}}_{\mathbf{w}\in\mathcal{W}}\langle\nabla\ell_{t}(\mathbf{w}_{t}),\mathbf{w}\rangle+\frac{1}{2\eta}\|\mathbf{w}-\mathbf{w}_{t}\|_{\tilde{H}_{t}}^{2}$. Then, we have
$\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\|_{H_{t}}\leq 2\eta\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{H_{t}^{-1}}\leq\frac{2\eta}{\lambda}\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{2}\leq\frac{4\eta}{\lambda}.$
Proof of Lemma F.9.
We make a slight improvement to Lemma 20 of Zhang and Sugiyama [51].
Lemma F.10 (Lemma 20 of Zhang and Sugiyama 51).
Under the same conditions as Lemma F.9, the following holds:
$\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\|_{H_{t}}\leq 2\eta\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{H_{t}^{-1}}\leq\frac{2\eta}{\lambda}\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{2}\leq\frac{2\eta\sqrt{K}}{\lambda}.$
In the proof of Lemma 20 by Zhang and Sugiyama [51], the bound for $\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{2}$ was given as $\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{2}\leq 2\sqrt{K}$. However, a tighter bound can be established: $\|\nabla\ell_{t}(\mathbf{w}_{t})\|_{2}\leq 2$. Using this tighter bound, we can refine and derive the improved result. ∎
@@ Appendix G Proofs of Theorem 3
In this section, we provide the proof of Theorem 3.
@@ G.1 Adversarial Rewards Construction
We follow a similar context construction as in Section D.1, but now consider the case where all context vectors are distinct and the rewards are non-uniform. Accordingly, the total number of items is $N=\binom{d}{d/4}$. Given $\mathbf{w}_{V}$, we define a unique item $i^{\star}\in[N]$ as the one that maximizes $x_{i}^{\top}\mathbf{w}_{V}$, i.e., $x_{i^{\star}}=x_{V}$. In addition, let $i^{\prime}\in[N]\setminus{i^{\star}}$ denote the item with the second-largest value of $x_{i}^{\top}\mathbf{w}_{V}$. At each round $t$, we assign a reward of $1$ to $i^{\star}$ when $t$ is odd, and a reward of $1$ to one randomly selected item other than $i^{\star}$ when $t$ is even. For simplicity, let ${\mathcal{T}}^{\star}\subseteq[T]$ denote the set of rounds corresponding to the former case, i.e., when the reward $1$ is assigned to $i^{\star}$. Specifically, for $t\in{\mathcal{T}}^{\star}$, the rewards are constructed as follows:
$r_{ti}=\begin{cases}1,&\text{for }i=i^{\star}\\ \nu,&\text{for }i=i^{\prime}\\ \gamma,&\text{otherwise},\end{cases}$
where $\nu$ and $\gamma$ are defined as
$\nu=\frac{e^{d/4-0.5}}{v_{0}+e^{d/4-0.5}},$ (G.1)
and
$\gamma=\min_{S\in\mathcal{S}}\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})\nu}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}=\frac{\nu}{v_{0}+1}<1.$
Moreover, for rounds $t\notin{\mathcal{T}}^{\star}$, let $\hat{i}\in[N]\setminus{i^{\star}}$ denote the randomly selected item (other than $i^{\star}$) that receives a reward of $1$. The rewards are then defined as follows:
$r_{ti}=\begin{cases}1,&\text{for }i=\hat{i}\\ \nu,&\text{for }i=i^{\prime}\\ \gamma,&\text{otherwise}.\end{cases}$
@@ G.2 Main Proof of Theorem 3
We begin the proof by introducing useful lemmas.
@@
Let $R_{t}(S^{\star}_{t},\mathbf{w}_{V})=\frac{\sum_{i\in S^{\star}_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})r_{ti}}{v_{0}+\sum_{j\in S^{\star}_{t}}\exp(x_{j}^{\top}\mathbf{w}_{V})}$. Then, we have
$S^{\star}_{t}=\begin{cases}\{i^{\star}\}&\text{for }t\in{\mathcal{T}}^{\star}\\ \{\hat{i},i^{\prime}\}&\text{for }t\notin{\mathcal{T}}^{\star}.\end{cases}$
Lemma G.1 establishes that any policy which relies solely on rewards—without estimating the underlying parameter—incurs a regret of order $\Omega(T)$ in this instance. In particular, the reward distributions over rounds in $\mathcal{T}$$^{\star}$ and $T]\setminus{\mathcal{T}}^{\star}$ are identical up to item permutations, making them indistinguishable based solely on the given rewards. Consequently, reward-based heuristics—such as selecting the top-1 or top-2 items, or choosing items whose rewards exceed a fixed threshold—necessarily suffer $\Omega(T)$ regret on $\mathcal{T}$$^{\star}$, $T]\setminus{\mathcal{T}}^{\star}$, or both. For instance, the top-1 strategy incurs zero regret for $t\in{\mathcal{T}}^{\star}$ but $\Omega(T)$ regret for $t\notin{\mathcal{T}}^{\star}$, since the optimal assortment size in that case is 2. Conversely, the top-2 strategy suffers $\Omega(T)$ regret for $t\in{\mathcal{T}}^{\star}$ and zero regret for $t\notin{\mathcal{T}}^{\star}$. Similarly, a threshold-based policy that selects items whose rewards exceed a fixed value yields a fixed-size assortment for all $t$ (since the reward distributions are identical across rounds), and therefore necessarily incurs $\Omega(T)$ regret in one or both cases. Therefore, we can safely exclude such reward-based heuristics from consideration and focus our analysis on the rounds in $\mathcal{T}$$^{\star}$.
@@
For any $t\in{\mathcal{T}}^{\star}$ and $i^{\star}\notin S_{t}$, we have
$R_{t}(S_{t},\mathbf{w}_{V})\leq\frac{\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}.$
Lemma G.2 shows that, for $t\in{\mathcal{T}}^{\star}$ and $i^{\star}\notin S_{t}$, the expected revenue of $S_{t}$ can be upper bounded by that of a singleton set whose reward has been replaced with $1$.
Now, we are ready to provide the proof of Theorem 3. Let $x_{U_{1}},\dots,x_{U_{L}}$ be the distinct feature vectors contained in assortments $S_{t}$ with $U_{1},\dots,U_{L}\in\mathcal{V}_{d/4}$. Let $\tilde{U}_{t}$ be the (unique) item among $U_{1},\dots,U_{L}$ that maximizes $x_{U}^{\top}\mathbf{w}_{V}$, i.e., $\tilde{U}_{t}=\mathop{\mathrm{argmax}}_{U\in\{U_{1},\dots,U_{L}\}}x_{U}^{\top}\mathbf{w}_{V}$, where $\mathbf{w}_{V}$ is the underlying parameter. Then, we have
$\sum_{t=1}^{T}$ $R_{t}(S^{\star}_{t},\mathbf{w}_{V})-R_{t}(S_{t},\mathbf{w}_{V})$
$\geq\sum_{t\in{\mathcal{T}}^{\star}}R_{t}(S^{\star}_{t},\mathbf{w}_{V})-R_{t}(S_{t},\mathbf{w}_{V})$
$=\sum_{t\in{\mathcal{T}}^{\star}}\frac{\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{V}^{\top}\mathbf{w}_{V})}-R_{t}(S_{t},\mathbf{w}_{V})$
$\geq\sum_{t\in{\mathcal{T}}^{\star}}\mathbf{1}\{i^{\star}\notin S_{t}\}\left(\frac{\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{V}^{\top}\mathbf{w}_{V})}-\frac{\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\max_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\right)$
$=\sum_{t\in{\mathcal{T}}^{\star}}\mathbf{1}\{i^{\star}\notin S_{t}\}\left(\frac{\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{V}^{\top}\mathbf{w}_{V})}-\frac{\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{\tilde{U}_{t}}^{\top}\mathbf{w}_{V})}\right),$
where the first equality holds by Lemma G.1 (and recall that $x_{i^{\star}}=x_{V}$), and the second inequality holds by Lemma G.2. Then, by applying Lemma D.1 with $K=1$, we can further bound the regret as follows:
$\sum_{t=1}^{T}R_{t}(S^{\star}_{t},\mathbf{w}_{V})-R_{t}(S_{t},\mathbf{w}_{V})\geq\frac{v_{0}}{(v_{0}+e)^{2}}\sum_{t\in{\mathcal{T}}^{\star}}\frac{\delta\epsilon}{2\sqrt{d}},$
where $\epsilon\in(0,1/d\sqrt{d})$ and $\delta=d/4-|\tilde{U}_{t}\cap V|$. Note that for $i^{\star}\in S_{t}$, we have $\tilde{U}_{t}=V$, and hence we get $\delta=0$.
The remainder of the analysis follows directly from the proof of Theorem 1 with $K=1$ and ${\mathcal{T}}^{\star}=T/2$. Therefore, setting $v_{0}=\Theta(1)$, we obtain
$\sup_{\mathbf{w}}\mathbb{E}_{\mathbf{w}}^{\pi}\left[\textup{{Reg}}_{T}(\mathbf{w})\right]$ $\geq\frac{1}{|\mathcal{V}_{d/4}|}\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{\mathbf{w}_{V}}^{\pi}\sum_{t=1}^{T}R_{t}(S^{\star}_{t},\mathbf{w}_{V})-R_{t}(S_{t},\mathbf{w}_{V})=\Omega\left(d\sqrt{T}\right).$
@@ G.3 Proofs of Lemmas for Theorem 3
First, for any $t\in[T]$ and all $i\in S^{\star}_{t}$, we show that $r_{ti}\geq R_{t}(S^{\star}_{t},\mathbf{w}_{V})$. We prove this by contradiction. Assume that there exists $i\in S^{\star}_{t}$ such that $r_{ti}<R_{t}(S^{\star}_{t},\mathbf{w}_{V})$. Then, removing the item $i$ from the assortment $S^{\star}_{t}$ yields higher expected revenue. This contradicts the optimality of $S^{\star}_{t}$. Thus, for any $t\in[T]$, we have
$r_{ti}\geq R_{t}(S^{\star}_{t},\mathbf{w}_{V}),\quad\forall i\in S^{\star}_{t}.$ (G.2)
Hence, for $t\in{\mathcal{T}}^{\star}$, we have $i^{\star}\in S_{t}^{\star}$, since $r_{i^{\star}}=1>R_{t}(S^{\star}_{t},\mathbf{w}_{V})$. Furthermore, for any other item $i\neq i^{\star}$, by definition of $\nu$ in Equation (G.1), we have
$r_{ti}\leq\nu=\frac{e^{d/4-0.5}}{v_{0}+e^{d/4-0.5}}<\frac{e^{d/4}}{v_{0}+e^{d/4}}=\frac{\exp(x_{V}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{V}^{\top}\mathbf{w}_{V})}=R_{t}(\{i^{\star}\},\mathbf{w}_{V}).$
Therefore, by Equation (G.2), only the item $i^{\star}$ is included in the optimal assortment for all $t\in{\mathcal{T}}^{\star}$.
We next analyze the case $t\notin{\mathcal{T}}^{\star}$. Using the same reasoning as above and applying Equation (G.2), we obtain that $\hat{i}\in S_{t}^{\star}$, since $r_{t\hat{i}}=1>R_{t}(S_{t}^{\star},\mathbf{w}_{V})$. Moreover, we have
$r_{ti^{\prime}}=\nu=\frac{e^{d/4-0.5}}{v_{0}+e^{d/4-0.5}}>\frac{e^{d/4-1}}{v_{0}+e^{d/4-1}}=R_{t}(\{\hat{i}\},\mathbf{w}_{V}),$
which implies $R_{t}(\{\hat{i}\},\mathbf{w}_{V})<R_{t}(\{\hat{i},i^{\prime}\},\mathbf{w}_{V})$. And for any $i\notin\{\hat{i},i^{\prime}\}$,
$r_{ti}=\gamma=\frac{\nu}{v_{0}+1}<\frac{1}{v_{0}+1}<R_{t}(\{\hat{i}\},\mathbf{w}_{V})<R_{t}(\{\hat{i},i^{\prime}\},\mathbf{w}_{V}).$
Therefore, by Equation (G.2), we can conclude that $S^{\star}_{t}=\{\hat{i},i^{\prime}\}$. This completes the proof. ∎
@@
Case 1. $i^{\prime}\in S_{t}$ and $i^{\star}\notin S_{t}$.
@@
$\gamma=\min_{S\in\mathcal{S}}\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})\nu}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})\nu}{v_{0}+\min_{i\in S_{t}}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})\nu}{v_{0}+\exp(x_{i^{\prime}}^{\top}\mathbf{w}_{V})}.$ (G.3)